The energy of hydrogen atom in its ground state is -13.6 ev.the energy of the level correspoding to n =5

Grade:

1 Answers

Nirmal Singh.
10 years ago
The Energy at nth Orbit is En = E1 / n2
Where E1 is the energy at ground state = -13.6 eV
The energy at 5th orbit i.e n = 5
E5 = -13.6 / 52 = 0.544 eV
Thanks & Regards,
Nirmal Singh
Askiitians faculty